OpenStudy (anonymous):

Does anyone get this?? Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

slope of the line=m=(y2-y1)/(x2-x1)=-1/7 equation of the line: y-y1=m(x-x1) y-1=(-1/7)(x-5) 7y-1=-x+5 x+7y=6

OpenStudy (anonymous):

thanks. would you know this one? Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5